Write the equation of the line with the given slope of -4 and x-intercept = 1.
Answer by mananth(16946)

m=-4, x intercept = 1
The co ordinates of the the point of intersection with the X axis is (1,0)


equation in slope point form
y-y1= m(x-x1)
y-0) = -4 ( x- (1))
y= -4x +4
4x +y =4